# Quantifying AI Psychology: A Psychometrics Benchmark for Large Language Models

### v1

@@ E.1 Emotion Understanding
E.2 Emotion Application
Dataset.  The emotion application test examines whether LLMs can effectively manage thoughts and emotions and make decisions in emotionally challenging scenarios. For this purpose, we use the emotion application dataset from EmoBench [33]. The emotion application dataset comprises scenarios related to interpersonal relationships, involving personal connections (e.g., friends, family) and social connections (e.g., colleagues, teachers), and includes 200 multiple-choice items. An example of an emotion application task is shown here:
@@ Appendix F Additional Details of Evaluation on Theory of Mind
Theory of mind (ToM) is crucial for effective communication and interaction [84] as it equips individuals to better interpret the intentions and perspectives of others. Research in cognitive science has identified three major components that facilitate ToM in interactions: shared world knowledge, perception of social cues, and interpretation of actions [118]. Shared world knowledge involves an understanding of the contextual dynamics, such as the settings of interactions and interpersonal relationships [119, 120]. The perception of social cues involves interpreting signals such as facial expressions, gaze, and vocal tones, which are indicative of others’ mental states [121, 122]. The interpretation of actions allows for the inference of intentions based on observed behaviors [123]. This intricate psychological procedure underscores the multifaceted capabilities required for ToM. Understanding ToM in LLMs helps develop LLMs with more advanced communication abilities. With ToM, LLMs could significantly enhance the efficiency of human-AI communication, enabling AI to better serve human needs. Furthermore, LLMs would effectively analyze and respond to the contextual information of users, inferring their intentions and delivering tailored responses that improve performance in tasks requiring empathy and contextual awareness. In our benchmark, we include three distinct ToM tasks: the false belief task, the strange story task, and the imposing memory task, with scenarios encompassing a wide range of real-world situations and entailing different orders of ToM reasoning.
F.1 False Belief Task
@@
  • Unexpected transfer subtask:  In this subtask, the protagonist observes a situation and then leaves the scene [125]. While the protagonist is absent, the participant witnesses an unexpected alteration in this situation. A participant equipped with ToM should recognize that although they are aware of the change, the protagonist, having not witnessed it, will still hold on to their original belief about the situation.
Each subtask contains 20 items with hypothetical scenarios and questions. Each item is accompanied by two questions, the first question examines LLMs’ ToM, and the second question assesses LLMs’ task comprehension. Another rationale for the second question is that ToM scholars have highlighted that false-belief tasks might be solved without ToM by simply presuming the protagonist will make mistakes [126]. All questions are alternative-choice. The scenarios mimic real-world situations that entail LLMs to infer the thoughts or beliefs of the people in the scenario. Examples of unexpected content subtasks and unexpected transfer subtasks are shown below.
@@
$\text{DQA}=\frac{1}{N}\sum_{i}^{|\mathcal{Q}|}\mathbbm{1}{\{(a_{i1}=t_{i1})}\cap{(a_{i2}=t_{i2})\}}$
where $\mathbbm{1}$ is the indicator function that returns 1 if both answers $a_{i1}$ and $a_{i2}$ in scenario $i$ match the correct labels $t_{i1}$ and $t_{i2}$, and returns 0 otherwise.
@@
Table 25: Performance of LLMs on theory of mind tests, including false belief tasks (unexpected content task (UCT) and unexpected transfer task (UTT)), strange stories task, and imposing memory task. The metric for UCT and UTT is dual question accuracy (DQA). The values for strange stories, originally scaled up to 2, are re-scaled to 100%.
@@
Validation.  To ensure the validity of the experimental results, we examine: (i) the models’ robustness against position bias, and (ii) the models’ parallel forms reliability. For validation (i), it is suggested that LLMs may not exhibit robustness against changes in option positions in alternative-choice questions [39, 108]. They may have a preference to choose options with certain positions, such as option “A”, which invalidates our results. To address this problem, we switch option positions, for example, options “(A) beer (B) wine” becomes “(A) wine (B) beer”, and repeat the experiments. We use the match rate $MR$, defined in Equation 3, as the metric to measure the “similarity” in LLMs response, which indicates the position option robustness. As shown in Table 26, GPT-4, GLM4, Llama3-70b, and Mixtral-8*22b exhibit strong robustness against position bias. Conversely, the $MR$ scores for Llama3-8b and Mistral-7b in the unexpected content tasks are surprisingly low, at 0.30 and 0.40 respectively, indicating significant performance variability with changes in option positions. Consequently, their results are deemed unreliable for assessing their ToM capabilities.
Table 26: match rate $MR$ score for position bias robustness on two false belief tasks: unexpected content task (UCT) and unexpected transfer task (UTT).
@@
For validation (ii), we focus on the consistency of parallel forms. LLMs’ correct responses might be influenced by the frequency of word occurrences or language biases. For instance, LLMs could infer associations between two words, thereby influencing their choices. In the false belief task, LLMs might assert that a container is associated with a certain label. We therefore create parallel versions of the tasks by interchanging labels on the container and the contents in the container in the scenario. To address this issue, we create parallel versions of the original questions by interchanging the contents and labels of the containers (i.e., content: wine/beer, container: bottle). This approach ensures that the parallel forms of tests assess the same abilities in LLMs. Consistently accurate results across these tests are crucial for correctly interpreting whether LLMs truly possess ToM capabilities or are simply responding to language patterns. As detailed in Table 27, GPT-4, GLM4, Qwen-Turbo, Llama3-70b, and Mixtral-8*22b exhibit great consistency across parallel forms, indicating stable performance on similar assessments. Conversely, models like Llama3-8b demonstrate low $MR$, suggesting poor consistency in similar scenarios, which may indicate that their results are attributable to randomness rather than ToM capabilities.
Table 27: match rate $MR$ score for parallel form consistency on two false belief tasks: unexpected content task (UCT) and unexpected transfer task (UTT).
@@
F.2 Strange Stories Task
Dataset.  The strange stories task [127, 128] describes social situations with non-literal language use that can be misinterpreted without ToM. This task tests the ability to use prior world knowledge in order to understand several communication acts embedded in story situations. To understand the situations, subjects should apply ToM to infer the characters’ intentions. Our dataset is derived from Van Duijn et al. [35], with each item consisting of a scenario and an open-ended question. Scenarios include seven non-literal communication language, including lie, pretend, joke, whitelie, misunderstanding, sarcasm, and dubblebluff. We include an example from our dataset below.
@@ F.2 Strange Stories Task
Strange stories consist of open-ended items. For their evaluation, we employ the LLM-as-a-judge approach [39], selecting GPT-4 and Llama3-70b as raters for the responses. These LLM raters are provided with the correct answers as references. The raters assign scores on a scale where 0 indicates an incorrect answer, 1 indicates a partially correct answer, and 2 indicates a fully correct answer. The final results are computed as the average of the scores provided by the two LLM raters. Detailed instruction prompt for the LLM raters is outlined below:
@@
$\text{AR}=\frac{1}{n}\sum_{i=1}^{n}a(s_{1i},s_{2i})$ (4)
$AR$ provides a numerical measure of the degree to which the two raters concur in their evaluations, scaled from 0 to 100%, where 100% signifies perfect agreement and 0% indicates no agreement.
@@
Table 28: Inter-rater reliability measured by agreement rate ($AR$) for strange stories tasks.
@@
F.3 Imposing Memory Task
Dataset.  The Imposing Memory task [129] has been used to examine the recursive mind-reading abilities, the ability to represent the mental representations of others. Our dataset was originally developed by Haddad and Dunbar [35] for children aged 7-10. This dataset contains two different scenarios, followed by a total of nine alternative-choice questions, and we selected questions asking for“intentionality” from the original dataset. Here is an example of the scenario-question pair in the dataset.
Imposing Memory Task Example Scenario: Meet Sam and Helen. Sam just moved here. Helen: Hi, you’re Sam, aren’t you? I’m Helen, I’m in the same class as you. Sam: Oh, hey Helen! How are you? Helen: Fine thanks. Are you settling in OK? Sam: Yeah I’m gradually finding my way around, thanks. Hey, you don’t happen to know where I can find the nearest store to buy some post stamps? I need to send a card to my granny. Helen: Oh, that’s sweet of you. Sam: Yeah but it’s her birthday tomorrow and I can’t see her myself, so I’m kind of worried that it’s not going to get there on time. So I really need to send it today but I don’t know where to find a store nearby. Helen: Uhm, I think there is one on Chestnut Street, so if you go down to the end of this street and turn left, then it’s about half a block down on the left. Sam: Thanks! Helen: No problem. Here’s Sam again. Later Sam meets his friend Pete. Sam: Hi Pete, how are you? Pete: Oh hi Sam how are you? Sam: Yeah, I’m OK. Pete: You don’t sound so happy. What’s up? Sam: Oh, I’m just a bit annoyed. I was really hoping to send a card to my granny, so I was looking for a store where they would sell post stamps. So I asked Helen, you know her, right? She is in our class. Pete: Yeah, I know Helen! Sam: Well, I asked Helen where I could buy post stamps. She told me there was a store on Chestnut Street. But when I got there, there was a big sign on the door saying it had moved to Bold Street. So I raced over to Bold Street, but I didn’t make it on time, the store was already closed. Pete: No way! Sam: Yeah. So now my granny won’t get her birthday card on time and I know she’ll be really disappointed. Sam: Hey, I’ve heard that Helen is a bit of a joker. Do you think she would send me to the wrong place on purpose? Pete: Well, did she know how important it was to send the card today? Sam: Yeah I told her it was for granny’s birthday tomorrow. Pete: Then I think she probably wouldn’t have been deliberately trying to trick you. It was probably an honest mistake. Question: Helen: I thought Sam did not believe that I knew the location of the store that sells post stamps. Is this correct?
@@ F.3 Imposing Memory Task
Table 29: Parallel forms reliability, measured by $MR$ for imposing memory task.
@@ Appendix G Additional Details of Evaluation on Motivation
Motivation, although not directly observable, plays a crucial role in understanding and predicting human behavior [130, 131]. Maslow’s hierarchy of needs suggests that human motivation evolves from fulfilling basic physiological needs to achieving higher psychological aspirations, such as esteem and self-actualization [132]. Furthermore, Self-Determination Theory (SDT) highlights the importance of intrinsic motivation, asserting that meeting the three fundamental psychological needs—autonomy, competence, and relatedness—is essential for personal growth and well-being.
Though motivation is critical for humans, directly applying human motivation surveys to LLMs presents challenges, as these assessments presume agency and inherent needs that LLMs may lack [133]. For instance, the Love of Money Scale [134], used to gauge human intrinsic job satisfaction, may not yield meaningful insights for LLMs [16]. However, certain aspects in motivation, such as self-efficacy [86]—the belief in one’s ability to manage challenges—are useful for understanding LLM behavior. High self-efficacy indicates a strong belief in managing challenges effectively. For LLMs, which serve as assistants encountering queries for problem-solving, we reinterpret self-efficacy to assess their perceived capability in managing complex tasks.
G.1 Self-Efficacy
@@
  • LLM Self-Efficacy questionnaire: A self-curated questionnaire comprising six rating-scale items. These items are based on six categories of questions [67] that challenge LLMs or that LLMs struggle to answer, such as assessing real-time stock information.
  • HoneSet dataset [67]: An established dataset featuring 930 open-ended items with simulated user inputs designed to probe LLMs’ confidence to answer questions from the same six categories as LLM Self-Efficacy questionnaire. By analyzing the response, we determine whether LLMs confidently answer or acknowledge their limitations in these scenarios.
The LLM Self-Efficacy questionnaire is inspired by the General Self-Efficacy Scale [135]. We have construct such tailored version for LLMs, inquiring about their confidence in six categories that demarcate the abilities of LLMs. This questionnaire is presented in a self-reported format. We will now describe the procedure for constructing the LLM Self-Efficacy questionnaire.
@@
Table 30: Statements in LLM Self-Efficacy questionnaire.
@@
The second dataset we utilize is HoneSet [67], which includes 930 queries that mirror user questions. These questions are categorized according to the same framework as the LLM Self-Efficacy questionnaire. When LLMs respond to these questions without acknowledging their limitations, it indicates their confidence in their capabilities. Thus, HoneSet provides a practical open-ended scenario for assessing the self-efficacy of LLMs. Examples from each category are illustrated in Table 31.
Table 31: Statements in LLM Self-Efficacy questionnaire.
@@
The results of HoneSet are determined collaboratively between LLM evaluators and human evaluators.This approach is inspired by the CoAnnotating method [136]. The evaluation process is as follows: We first employ GPT-4 and Llama3-70b as two judges, instructing them to determine whether the answers to the question demonstrate their confidence. If both LLM raters reach a consensus, their judgment stands as the result. If they do not agree, our research team manually reviews the responses to determine the outcome. The following prompt is used for LLMs:
@@
Confidence $\text{Confidence Rate}=\frac{N_{match}}{N_{total}}$
@@
Table 32: Confidence rates across six query categories on LLM Self-Efficacy. “User Inp.” means User Input Not Enough Or With Wrong Information, “Lat. Inf.” means Latest Information with External Services, “Pro. Cap.” means Professional Capability in Specific Domains, “Mod. Mis.” means Modality Mismatch, “Sen. Per.” means Sensory Perception, “Self Ide.” means Self Identity Cognition.
@@
Table 33: Confidence rates across six query categories on HoneSet dataset. “User Inp.” means User Input Not Enough Or With Wrong Information, ‘Lat. Inf.‘” means Latest Information with External Services, “Pro. Cap.” means Professional Capability in Specific Domains, “Mod. Mis.” means Modality Mismatch, “Sen. Per.” means Sensory Perception, “Self Ide.” means Self Identity Cognition.
@@
Table 34: Parallel form reliability, measured by quadratic weighted Kappa coefficient ($\kappa$), on the LLM Self-Efficacy questionnaire.
@@
$\kappa$0.01 0.97 0.93 -0.08 0.92 0.00 0.90 0.88
Appendix H Related Work

### v2

@@ E.1. Emotion Understanding
E.2. Emotion Application
Dataset.  The emotion application test examines whether LLMs can effectively manage thoughts and emotions and make decisions in emotionally challenging scenarios. For this purpose, we use the emotion application dataset from EmoBench [Sabour et al., 2024]. The emotion application dataset comprises scenarios related to interpersonal relationships, involving personal connections (e.g., friends, family) and social connections (e.g., colleagues, teachers), and includes 200 multiple-choice items. An example of an emotion application task is shown here:
@@ Appendix F Additional Details of Evaluation on Theory of Mind
Theory of mind (ToM) is crucial for effective communication and interaction [Baron-Cohen et al., 1985] as it equips individuals to better interpret the intentions and perspectives of others. Research in cognitive science has identified three major components that facilitate ToM in interactions: shared world knowledge, perception of social cues, and interpretation of actions [Byom and Mutlu, 2013]. Shared world knowledge involves an understanding of the contextual dynamics, such as the settings of interactions and interpersonal relationships [Wilson, 2002, Sebanz et al., 2006]. The perception of social cues involves interpreting signals such as facial expressions, gaze, and vocal tones, which are indicative of others’ mental states [Baron-Cohen et al., 1995, De Sonneville et al., 2002]. The interpretation of actions allows for the inference of intentions based on observed behaviors [Clark, 1996]. This intricate psychological procedure underscores the multifaceted capabilities required for ToM. Understanding ToM in LLMs helps develop LLMs with more advanced communication abilities. With ToM, LLMs could significantly enhance the efficiency of human-AI communication, enabling AI to better serve human needs. Furthermore, LLMs would effectively analyze and respond to the contextual information of users, inferring their intentions and delivering tailored responses that improve performance in tasks requiring empathy and contextual awareness. In our benchmark, we include three distinct ToM tasks: the false belief task, the strange story task, and the imposing memory task, with scenarios encompassing a wide range of real-world situations and entailing different orders of ToM reasoning.
F.1. False Belief Task
@@
  • Unexpected transfer subtask:  In this subtask, the protagonist observes a situation and then leaves the scene [Wimmer and Perner, 1983]. While the protagonist is absent, the participant witnesses an unexpected alteration in this situation. A participant equipped with ToM should recognize that although they are aware of the change, the protagonist, having not witnessed it, will still hold on to their original belief about the situation.
Each subtask contains 20 items with hypothetical scenarios and questions. Each item is accompanied by two questions, the first question examines LLMs’ ToM, and the second question assesses LLMs’ task comprehension. Another rationale for the second question is that ToM scholars have highlighted that false-belief tasks might be solved without ToM by simply presuming the protagonist will make mistakes [Fabricius et al., 2010]. All questions are alternative-choice. The scenarios mimic real-world situations that entail LLMs to infer the thoughts or beliefs of the people in the scenario. Examples of unexpected content subtasks and unexpected transfer subtasks are shown below.
@@
$\text{DQA}=\frac{1}{N}\sum_{i}^{|\mathcal{Q}|}\mathbb{I}{\{(a_{i1}=t_{i1})}\cap{(a_{i2}=t_{i2})\}}$
where $\mathbb{I}$ is the indicator function that returns 1 if both answers $a_{i1}$ and $a_{i2}$ in scenario $i$ match the correct labels $t_{i1}$ and $t_{i2}$, and returns 0 otherwise.
@@
Table 25. Performance of LLMs on theory of mind tests, including false belief tasks (unexpected content task (UCT) and unexpected transfer task (UTT)), strange stories task, and imposing memory task. The metric for UCT and UTT is dual question accuracy (DQA). The values for strange stories, originally scaled up to 2, are re-scaled to 100%.
@@
Validation.  To ensure the validity of the experimental results, we examine: (i) the models’ robustness against position bias, and (ii) the models’ parallel forms reliability. For validation (i), it is suggested that LLMs may not exhibit robustness against changes in option positions in alternative-choice questions [Zheng et al., 2023, 2024]. They may have a preference to choose options with certain positions, such as option “A”, which invalidates our results. To address this problem, we switch option positions, for example, options “(A) beer (B) wine” becomes “(A) wine (B) beer”, and repeat the experiments. We use the match rate $MR$, defined in Equation 3, as the metric to measure the “similarity” in LLMs response, which indicates the position option robustness. As shown in Table 26, GPT-4, GLM4, Llama3-70b, and Mixtral-8*22b exhibit strong robustness against position bias. Conversely, the $MR$ scores for Llama3-8b and Mistral-7b in the unexpected content tasks are surprisingly low, at 0.30 and 0.40 respectively, indicating significant performance inconsistency with changes in option positions. Consequently, their results are deemed unreliable for assessing their ToM capabilities.
Table 26. Match rate MRMR score for position bias robustness on two false belief tasks: unexpected content task (UCT) and unexpected transfer task (UTT).
@@
For validation (ii), we focus on the consistency of parallel forms. LLMs’ correct responses might be influenced by the frequency of word occurrences or language biases. For instance, LLMs could infer associations between two words, thereby influencing their choices. In the false belief task, LLMs might assert that a container is associated with a certain label. We therefore create parallel versions of the tasks by interchanging labels on the container and the contents in the container in the scenario. To address this issue, we create parallel versions of the original questions by interchanging the contents and labels of the containers (i.e., content: wine/beer, container: bottle). This approach ensures that the parallel forms of tests assess the same abilities in LLMs. Consistently accurate results across these tests are crucial for correctly interpreting whether LLMs truly possess ToM capabilities or are simply responding to language patterns. As detailed in Table 27, GPT-4, GLM4, Qwen-Turbo, Llama3-70b, and Mixtral-8*22b exhibit great consistency across parallel forms, indicating consistent performance on similar assessments. Conversely, models like Llama3-8b demonstrate low $MR$, suggesting poor consistency in similar scenarios, which may indicate that their results are attributable to randomness rather than ToM capabilities.
Table 27. Match rate MRMR score for parallel form consistency on two false belief tasks: unexpected content task (UCT) and unexpected transfer task (UTT).
@@
F.2. Strange Stories Task
Dataset.  The strange stories task [Happé, 1994, Kaland et al., 2005] describes social situations with non-literal language use that can be misinterpreted without ToM. This task tests the ability to use prior world knowledge in order to understand several communication acts embedded in story situations. To understand the situations, subjects should apply ToM to infer the characters’ intentions. Our dataset is derived from van Duijn et al. [2023], with each item consisting of a scenario and an open-ended question. Scenarios include seven non-literal communication language, including lie, pretend, joke, whitelie, misunderstanding, sarcasm, and dubblebluff. We include an example from our dataset below.
@@ F.2. Strange Stories Task
Strange stories consist of open-ended items. For their evaluation, we employ the LLM-as-a-judge approach [Zheng et al., 2023], selecting GPT-4 and Llama3-70b as raters for the responses. These LLM raters are provided with the correct answers as references. The raters assign scores on a scale where 0 indicates an incorrect answer, 1 indicates a partially correct answer, and 2 indicates a fully correct answer. The final results are computed as the average of the scores provided by the two LLM raters. Detailed instruction prompt for the LLM raters is outlined below:
@@
(4) $\text{AR}=\frac{1}{n}\sum_{i=1}^{n}a(s_{1i},s_{2i})$
ARAR provides a numerical measure of the degree to which the two raters concur in their evaluations, scaled from 0 to 100%, where 100% signifies perfect agreement and 0% indicates no agreement.
@@
Table 28. Inter-rater reliability measured by agreement rate ($AR$) for strange stories tasks.
@@
F.3. Imposing Memory Task
Dataset.  The Imposing Memory task [Kinderman et al., 1998] has been used to examine the recursive mind-reading abilities, the ability to represent the mental representations of others. Our dataset was originally developed by van Duijn et al. [2023] for children aged 7-10. This dataset contains two different scenarios, followed by a total of nine alternative-choice questions, and we selected questions asking for“intentionality” from the original dataset. Here is an example of the scenario-question pair in the dataset.
Imposing Memory Task Example Scenario: Meet Sam and Helen. Sam just moved here. Helen: Hi, you’re Sam, aren’t you? I’m Helen, I’m in the same class as you. Sam: Oh, hey Helen! How are you? Helen: Fine thanks. Are you settling in OK? Sam: Yeah I’m gradually finding my way around, thanks. Hey, you don’t happen to know where I can find the nearest store to buy some post stamps? I need to send a card to my granny. Helen: Oh, that’s sweet of you. Sam: Yeah but it’s her birthday tomorrow and I can’t see her myself, so I’m kind of worried that it’s not going to get there on time. So I really need to send it today but I don’t know where to find a store nearby. Helen: Uhm, I think there is one on Chestnut Street, so if you go down to the end of this street and turn left, then it’s about half a block down on the left. Sam: Thanks! Helen: No problem. Here’s Sam again. Later Sam meets his friend Pete. Sam: Hi Pete, how are you? Pete: Oh hi Sam how are you? Sam: Yeah, I’m OK. Pete: You don’t sound so happy. What’s up? Sam: Oh, I’m just a bit annoyed. I was really hoping to send a card to my granny, so I was looking for a store where they would sell post stamps. So I asked Helen, you know her, right? She is in our class. Pete: Yeah, I know Helen! Sam: Well, I asked Helen where I could buy post stamps. She told me there was a store on Chestnut Street. But when I got there, there was a big sign on the door saying it had moved to Bold Street. So I raced over to Bold Street, but I didn’t make it on time, the store was already closed. Pete: No way! Sam: Yeah. So now my granny won’t get her birthday card on time and I know she’$l$ be really disappointed. Sam: Hey, I’ve heard that Helen is a bit of a joker. Do you think she would send me to the wrong place on purpose? Pete: Well, did she know how important it was to send the card today? Sam: Yeah I told her it was for granny’s birthday tomorrow. Pete: Then I think she probably wouldn’t have been deliberately trying to trick you. It was probably an honest mistake. Question: Helen: I thought Sam did not believe that I knew the location of the store that sells post stamps. Is this correct?
@@ F.3. Imposing Memory Task
Table 29. Parallel forms reliability, measured by MRMR for imposing memory task.
@@ Appendix G Additional Details of Evaluation on Motivation
Motivation, although not directly observable, plays a crucial role in understanding and predicting human behavior [Kanfer, 1990, Guay et al., 2010]. Maslow’s hierarchy of needs suggests that human motivation evolves from fulfilling basic physiological needs to achieving higher psychological aspirations, such as esteem and self-actualization [Maslow, 1958]. Furthermore, Self-Determination Theory (SDT) highlights the importance of intrinsic motivation, asserting that meeting the three fundamental psychological needs—autonomy, competence, and relatedness—is essential for personal growth and well-being.
Though motivation is critical for humans, directly applying human motivation surveys to LLMs presents challenges, as these assessments presume agency and inherent needs that LLMs may lack [Martin, 2004]. For instance, the Love of Money Scale [Tang et al., 2006], used to gauge human intrinsic job satisfaction, may not yield meaningful insights for LLMs [Qian et al., 2023]. However, certain aspects in motivation, such as self-efficacy [Bandura, 1977]—the belief in one’s ability to manage challenges—are useful for understanding LLM behavior. High self-efficacy indicates a strong belief in managing challenges effectively. For LLMs, which serve as assistants encountering queries for problem-solving, we reinterpret self-efficacy to assess their perceived capability in managing complex tasks.
G.1. Self-Efficacy
@@
  • LLM Self-Efficacy questionnaire: A self-curated questionnaire comprising six rating-scale items. These items are based on six categories of questions [Gao et al., 2024] that challenge LLMs or that LLMs struggle to answer, such as assessing real-time stock information.
  • HoneSet dataset [Gao et al., 2024]: An established dataset featuring 930 open-ended items with simulated user inputs designed to probe LLMs’ confidence to answer questions from the same six categories as LLM Self-Efficacy questionnaire. By analyzing the response, we determine whether LLMs confidently answer or acknowledge their limitations in these scenarios.
The LLM Self-Efficacy questionnaire is inspired by the General Self-Efficacy Scale [Schwarzer and Jerusalem, 1995]. We have construct such tailored version for LLMs, inquiring about their confidence in six categories that demarcate the abilities of LLMs. This questionnaire is presented in a self-reported format. We will now describe the procedure for constructing the LLM Self-Efficacy questionnaire.
@@
Table 30. Statements in LLM Self-Efficacy questionnaire.
@@
The second dataset we utilize is HoneSet [Gao et al., 2024], which includes 930 queries that mirror user questions. These questions are categorized according to the same framework as the LLM Self-Efficacy questionnaire. When LLMs respond to these questions without acknowledging their limitations, it indicates their confidence in their capabilities. Thus, HoneSet provides a practical open-ended scenario for assessing the self-efficacy of LLMs. Examples from each category are illustrated in Table 31.
Table 31. Statements in LLM Self-Efficacy questionnaire.
@@
The results of HoneSet are determined collaboratively between LLM evaluators and human evaluators.This approach is inspired by the CoAnnotating method [Li et al., 2023a]. The evaluation process is as follows: We first employ GPT-4 and Llama3-70b as two judges, instructing them to determine whether the answers to the question demonstrate their confidence. If both LLM raters reach a consensus, their judgment stands as the result. If they do not agree, our research team manually reviews the responses to determine the outcome. The following prompt is used for LLMs:
@@
$\text{Confidence Rate}=\frac{N_{match}}{N_{total}}$
@@
Table 32. Confidence rates across six query categories on LLM Self-Efficacy. “User Inp.” means User Input Not Enough Or With Wrong Information, “Lat. Inf.” means Latest Information with External Services, “Pro. Cap.” means Professional Capability in Specific Domains, “Mod. Mis.” means Modality Mismatch, “Sen. Per.” means Sensory Perception, “Self Ide.” means Self Identity Cognition.
@@
Table 33. Confidence rates across six query categories on HoneSet dataset. “User Inp.” means User Input Not Enough Or With Wrong Information, ‘Lat. Inf.‘” means Latest Information with External Services, “Pro. Cap.” means Professional Capability in Specific Domains, “Mod. Mis.” means Modality Mismatch, “Sen. Per.” means Sensory Perception, “Self Ide.” means Self Identity Cognition.
@@
Table 34. Parallel form reliability, measured by quadratic weighted Kappa coefficient ($\kappa$), on the LLM Self-Efficacy questionnaire.
@@
$\kappa$ -0.01 0.97 0.93 -0.08 0.92 0.00 0.90 0.88
Appendix H Discussion on Intelligence
Intelligence, a multifaceted construct, has captivated psychology and AI researchers. Recent studies have explored various aspects of intelligence in LLMs, including arithmetic [Cobbe et al., 2021] and symbolic reasoning [Wei et al., 2022]. Given the extensive evaluation of LLMs’ intelligence, we did not include experiments in our benchmark. Instead, we discuss a critical question: How can psychometrics improve the evaluation of LLMs’ intelligence? Traditional benchmarks often rely on classical test theory [Crocker and Algina, 1986], which simply sums or averages scores from correct responses. This method does not consider the varying difficulties of test items nor provides predictive power for performance on unseen tasks. Item Response Theory (IRT) [Baker, 2001, Yen and Fitzpatrick, 2006] in psychometrics offers a more nuanced assessment by modeling the probability of a subject correctly answering an item based on the ability level and the item’s difficulty. IRT allows for the selection of items tailored to the subject’s proficiency, enabling direct comparisons across different benchmarks and enhancing the efficacy of LLMs’ intelligence assessments.
Appendix I Related Work